$ 50,230 32,210 Prepaid expenses 20,000 5,912 1,560 The company also reported total assets of $353,816 thousand, total liabilities of $81,630 thousand, and a return on total assets of 18.1 percent Required: Round ratios to 1 decimal place to 1 < Prev 30f 4
Answer #1
a-1 (quick asset) a-2   (current asset) a-3 (current liabilities)
cash 50230 50230
marketable securities 55926 55926
Accounts receivable 23553 23553
Inventories 32210
Prepaid expense 5736
Note payable (due within a year) 20000
Accounts payable 5912
Dividend payable 1560
Accrued liabilities 21532
Income tax payable 6438
Total 129709 167655

55442

**Retained earning is part of equity

b)

Ratio Formula
Quick ratio quick asset /current liabilities 129709/55442 =2.3
Currentratio current asset /current liabilities 167655/55442 = 3.0
working capital current asset -current liabilities 167655-55442= 112213
Debt ratio Total liabilities /Total asset 81630/353816= 23.1%